OpenStudy (tardis):

The question is True or False: Does cos(x+y) = cos(x)+cos(y)? I'm having trouble figuring out how to justify my answer algebraically. Any help would be much appreciated and given a medal :)

OpenStudy (hezekieli):

Find a contradicting case to disprove it at least. Still figuring out it myself... :P

OpenStudy (anonymous):

if you need help visualizing it, try plugging in numbers. let's say x=30 degrees and y=60 degrees. would cos(90)=cos(30)+cos(60)? That gives you a starting point because you know if you should prove or disprove that statement.

zepdrix (zepdrix):

Use the Cosine Angle Sum Formula to justify your choice :)\[\large\rm \cos(A+B)=\cos A \cos B-\sin A \sin B\]Unless that's not what you meant by algebraically :3

OpenStudy (jojokiw3):

\[\cos(44) + \cos(46) \not = \cos(90)\]